OpenStudy (anonymous):

The value of x in the proportion 1/2:2/3 = 3/4:x is A.4/9 B.1 7/9 C.1/4 D. 1

OpenStudy (badhi):

It can also be written as,$$\frac{1}{2}\div\frac{2}{3}=\frac{3}{4}\div x$$ now find x

OpenStudy (anonymous):

Would it be 1?

OpenStudy (anonymous):

yea i think its 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

do i get a medal